OpenStudy (anonymous):

during the first part of a trip a canoist travels 34 miles at a certain speed. The canoeist travels 22 miles on the second part at a speed of 5 miles slower. the total time for the trip is 3 hrs what was the speed on each part of the trip?

OpenStudy (anonymous):

d = 34 + 22 = 55miles t = 3 hours v(average) = 55/3 mph v(average) = (v1 + (v1 - 5))/3 = 55/3 solve for v1 to get the speed on the first part of the trip, v2 = v1-5

OpenStudy (dumbcow):

Let x be the speed on 1st part of trip, t1 and t2 denote time for each trip distance = rate*time Set up equations: (i) 34 = x*t1 (ii) 22 = (x-5)*t2 (iii) t1 + t2 = 3 Goal is to solve for x, use substitution to get an equation in terms of only x solve (iii) for t1 t1 = 3 - t2 substitute it into (i) then solve for t2 34 = x(3-t2) 34 = 3x - x*t2 x*t2 = 3x -34 t2 = (3x -34)/x ---------------- Solve (ii) for t2 t2 = 22/(x-5) ---------------- Now set t2 = t2 in terms of x (3x-34)/x = 22/(x-5) Cross-multiply to get rid of fractions (x-5)(3x-34) = 22x 3x^2 -49x + 170 = 22x 3x^2 -71x + 170 = 0 Use quadratic formula: x = [71 +-sqrt(71^2 - 4*3*170)]/6 x = (71 +- 54.781)/6 x = 2.7 or 20.964 Since x is the faster speed and speed must be greater than 0, then (x-5)>0 so x=2.7 is irrelevant The speeds on each trip were 20.964 and 15.964